# Improved Stochastic Optimization of LogSumExp

### v2

@@ Abstract
The LogSumExp function, also known as the free energy, plays a central role in many important optimization problems, including entropy-regularized optimal transport and distributionally robust optimization (DRO). It is also the dual to the Kullback-Leibler (KL) divergence, which is widely used in machine learning. In practice, when the number of exponential terms inside the logarithm is large or infinite, optimization becomes challenging since computing the gradient requires differentiating every term. Previous approaches that replace the full sum with a small batch introduce significant bias. We propose a novel approximation to LogSumExp that can be efficiently optimized using stochastic gradient methods. This approximation is rooted in a sound modification of the KL divergence in the dual, resulting in a new $f$-divergence called the safe KL divergence. The accuracy of the approximation is controlled by a tunable parameter and can be made arbitrarily small. Like the LogSumExp, our approximation preserves convexity. Moreover, when applied to an $L$-smooth function bounded from below, the smoothness constant of the resulting objective scales linearly with $L$. Experiments in DRO and continuous optimal transport demonstrate the advantages of our approach over state-of-the-art baselines and the effective treatment of numerical issues associated with the standard LogSumExp and KL.
@@ 1 Introduction
$F(\varphi;\mu)\coloneqq\ln\int e^{\varphi(x)}\,\mathrm{d}\mu(x)\in(-\infty,\infty]$ (1)
@@
LogSumExp function appears commonly in optimization objectives, e.g., multiclass classification with softmax probabilities (bishop2006pattern), semi-dual formulation of entropy-regularized optimal transport (OT) (peyre2018computational; genevay2016stochastic), minimax problems (pee2011solving), distributionally robust optimization (DRO) (huKullbackLeiblerDivergenceConstrained; ben-tal_robust_2013; kuhnDistributionallyRobustOptimization2024), maximum likelihood estimation (MLE) for exponential families and graphical models (wainwright2008graphical), variational Bayesian methods (khan2018fast; khanBayesianLearningRule2023), information geometry (amariMethodsInformationGeometry2000), KL-regularized Markov decision processes (tiapkin2024demonstrationregularized). These problems involve minimizing $F(\varphi;\mu)$ w.r.t. a function $\varphi$, potentially parameterized by a vector $\theta$, e.g., a vector of neural network weights. Such optimization is characterized by two challenges. First, the decision variable $\varphi$ or $\theta$ often has large or infinite dimension. Second, the support of the measure $\mu$ can also be large or infinite. The first challenge is usually addressed by the use of first-order methods like Stochastic Gradient Descent (SGD), which are suitable for high-dimensional problems due to their cheap iterations. Unfortunately, there is no universal way to tackle the second challenge. If the number of exponential terms under the logarithm is large or infinite, i.e., the Monte Carlo estimate $\log\sum_{i=1}^{N}e^{\varphi(x_{i})}$ for a large $N$, the gradient computation requires differentiating each term, and, to the best of our knowledge, no cheap unbiased stochastic gradient have been proposed.
@@
In the current work, we propose a general-purpose approach to tackle both mentioned challenges. To that end, we use a SoftPlus approximation of $F(\varphi;\mu)$ that allows using stochastic gradient methods while remaining close to the original objective. We start with a variational formulation analogous to the one in the Gibbs principle, but with the KL-divergence replaced with another $f$-divergence – the safe KL divergence. The corresponding variational problem can be of interest itself, as it possesses some properties which can be beneficial compared to the KL penalty – e.g., uniform density bound. Moreover, it can be also viewed as an approximation of a conditional value at risk functional (CVaR). In fact, the same functional (with different parameters) appeared in soma2020statistical in the context of smooth CVaR approximation. Thus, we demonstrate that it generates a family of problems including CVaR and LogSumExp minimization as limit cases.
@@
Related works. We are not aware of any universal way to efficiently treat optimization objectives involving the LogSumExp functional (1), especially in infinite-dimensional settings. This functional appeared in different applications and was treated on a case-by-case basis. bouchard2007efficient studies three upper bounds on LogSumExp for approximate Bayesian inference. One of them is a particular case of the approximation proposed in the present work. aueb2016one constructs a bound on softmax probabilities and shows that it leads to a bound on LogSumExp in the context of multiclass classification. nielsen2016guaranteed approximate LogSumExp in the context of estimating divergences between mixture models. Their approximation combines LogSumExp bounds based on $\min$ and $\max$. Tucker2017REBAR; Luo2020SUMO propose and study unbiased estimators for latent variable models based on Russian Roulette truncation. LyneGirolamiEtAl2015RussianRoulette; SpringShrivastava2017LSHPartition focus on estimating the partition function itself, and do not consider questions of optimization involving the partition function. huKullbackLeiblerDivergenceConstrained and, susequently, levy2020large study DRO problems with $f$-divergences. They propose a batch-based approximation. When the ambiguity set is the unit simplex, and KL divergence penalty is used, the original objective is the LogSumExp of the losses over the entire dataset, while the approximation replaces it with the average of LogSumExp terms computed on individual batches. This approximation introduces bias, which can only be reduced by using large batch sizes. Deterministic LogSumExp maximization and minimization were considered in Selvi2020ConvexMaximization and KanNagyRuthotto2023LSEMINK, respectively. For stabilizing numerics related to evaluation of LogSumExp function, we refer to BlanchardHighamHigham2020AccurateLogSumExp; Higham2021WhatIsLogSumExp.
@@
    We empirically demonstrate the effectiveness of our approach on tasks, including computing continuous entropy-regularized OT and various DRO formulations [^1] The source code is available at https://github.com/egorgladin/logsumexp-approx.. Our method outperforms existing state-of-the-art baselines in these applications and circumvents the overflow issue (Remark 3.1). It can also be combined with existing techniques. Therefore, it serves as a versatile tool for solving large-scale optimization problems.
@@
Notation. Given $a,a_{1},\dots,a_{n}\in\mathbb{R}$, we define $\operatorname{LogSumExp}(a_{1},\dots,a_{n})\coloneqq\log(\sum_{i=1}^{n}e^{a_{i}})$ and $\operatorname{SoftPlus}(a)\coloneqq\log(1+e^{a})$. Given a measurable space $\mathcal{X}$, by $\mathcal{P}(\mathcal{X})$ we denote the space of probability measures on $\mathcal{X}$, and by $\mathcal{C}(\mathcal{X})$ the space of continuous functions on $\mathcal{X}$. Let $\mu,\nu\in\mathcal{P}(\mathcal{X})$. Define Kullback–Leibler (KL) divergence as
@@ 2 SoftPlus approximation of log-partition function
$F(\varphi;\mu)={\sup_{\nu}}\left\{\int_{{\mathcal{X}}}\varphi(x)\,\mathrm{d}\nu(x)-D_{KL}(\nu,\mu):\nu\in\mathcal{P}(\mathcal{X}),\,\int_{{\mathcal{X}}}|\varphi(x)|\,\mathrm{d}\nu(x)<\infty\right\}$ (2)
with the maximum attained at the Gibbs measure $d\nu^{*}(x)=e^{\varphi(x)-F(\varphi;\mu)}d\mu(x),$ once $F(\varphi;\mu)<\infty$, see (gibbs1902elementary, Chapter XI, Theorem VI) or (polyanskiy2025information, Proposition 4.7) for the modern treatment.
@@
It is easy to see that $f_{\rho}(t)\to f_{0}(t)\coloneqq t\log t+1-t$ as $\rho\to 0$. Since $f_{0}$ induces the standard KL-divergence, $D_{\rho}$ is its approximation with accuracy regulated by the parameter $\rho$.
@@
$F_{\rho}(\varphi;\mu)\coloneqq\sup_{{\nu}}\left\{\int_{{\mathcal{X}}}\varphi(x)\,\mathrm{d}\nu(x)-D_{\rho}(\nu,\mu):\right$.
$\left.\nu\in\mathcal{P}(\mathcal{X}),\,\int_{{\mathcal{X}}}|\varphi(x)|\,\mathrm{d}\nu(x)<\infty\right\}.$ (5)
(i.e., $F_{\rho}(\cdot;\mu)$ is the convex conjugate of $D_{\rho}(\cdot,\mu)$). Note that the last term in $f_{\rho}$ prevents the density $\frac{d\nu}{d\mu}$ from being too large. In particular, it can not be greater than $\frac{1}{\rho}$. This can make the safe KL divergence a reasonable choice for unbalanced OT or DRO, as it imposes a hard constraint on the reweighting unlike the standard $D_{KL}$. Moreover, it can also be used instead of the entropy penalization in regularized OT (cf. capacity constrained transport in (benamou2015iterative, section 5.2)).
Again, by the convex duality and the variational principle (see birrell2022divergence, Theorem 6), we state the following properties.
@@
    If $\varphi(x)\leq M$ for all $x\in\mathcal{X}$, then $F_{\rho}(\varphi;\mu)\geq F(\varphi;\mu)-\rho e^{M-F(\varphi;\mu)}$ for $\rho\in\left(0,e^{F(\varphi;\mu)-M}\right)$.
@@
Let $a_{1},\dots,a_{n}\in\mathbb{R}$. Then for any $0<\rho<1$
$\operatorname{LogSumExp}(a_{1},\dots,a_{n})-\rho\leq\inf_{\alpha\in\mathbb{R}}\alpha-1+\frac{1}{\rho}\sum_{i=1}^{n}\log(1+\rho e^{a_{i}-\alpha})\leq\operatorname{LogSumExp}(a_{1},\dots,a_{n}).$
For $\rho=1$ our approximation coincides with Bouchard’s bound for LogSumExp (bouchard2007efficient).
@@ 2.1 Links to CVaR
$\operatorname{CVaR}_{\rho}(\varphi;\mu)\coloneqq\mathbb{E}_{X\sim\mu}\left[\varphi(X)\middle|\varphi(X)\geq Q_{1-\rho}\right]=\frac{1}{\rho}\int_{\varphi(x)\geq Q_{1-\rho}}\varphi(x)\,\mathrm{d}\mu(x),$
where $Q_{1-\rho}$ is the $(1-\rho)$-quantile of $\varphi(X)$, $X\sim\mu$ (rockafellar2000optimization). Moreover, by Theorem 1 in rockafellar2000optimization CVaR also has the following variational formulation:
@@
Remarkably, in soma2020statistical the authors obtained a smooth approximation to CVaR which, up to an additive constant, has the same form as $F_{\rho}$. However, they considered the approximation w.r.t. a different parameter—a ”temperature” inside SoftPlus. Finally, levy2020large proposed another similar smoothed version of CVaR (KL-regularized CVaR) in the context of DRO. For our approximation, we obtain the following bounds.
@@
$\operatorname{CVaR}_{\rho}(\varphi;\mu)+\lambda(\log\rho-1)\leq\lambda F_{\rho}(\varphi/\lambda;\mu)\leq\operatorname{CVaR}_{\rho}(\varphi;\mu)+\lambda\left(\log\rho-1+\frac{1}{\rho}\right).$ (10)
@@ 2.2 The case of parametric models
In some applications, the function $\varphi$ is defined as the parametric loss function $L(x,\theta)$ and the goal is to minimize objective involving (1) w.r.t. parameter $\theta$ to find the best model from the parametric family. In this section, we study our approximation to (1) in this parametric setting. Fix some closed parameter set $\Theta\subset\mathbb{R}^{d}$ and a loss function $L\colon\mathcal{X}\times\Theta\to\mathbb{R}$. Combining our approximation (6) and the minimization w.r.t. parameter $\theta$, we obtain the following minimization problem (note that we shifted $\alpha$ by $\log\rho$ compared to (6))
$\min_{\theta\in\Theta,\alpha\in\mathbb{R}}G_{\rho}(\theta,\alpha)\coloneqq\alpha+\log\rho-1+\frac{1}{\rho}\int_{{\mathcal{X}}}\log\left(1+e^{L(x,\theta)-\alpha}\right)\,\mathrm{d}\mu(x).$
@@
Note that $f_{\rho}(t)=\frac{1}{\rho}\left((\rho t)\log(\rho t)+(1-\rho t)\log(1-\rho t)\right)+1-t\log\rho.$ Thus, unlike the KL entropy function $t\log t+1-t$, $f_{\rho}$ possesses the following favorable properties:
@@
The above properties are important from the computational optimization point of view. Recall that $\frac{d}{dt}\log(1+e^{t})=\frac{e^{t}}{1+e^{t}}\eqqcolon\sigma(t)$. Thus, we immediately obtain the following formulas for the gradient:
@@ 3 Applications
In this section we consider several particular applications involving the objective (1) and show numerically, that our general-purpose approach based on approximation (6) leads to better performance of SGD-type algorithms than the baseline algorithms designed specifically for these applications. Source code for all experiments is available at https://github.com/egorgladin/logsumexp-approx.

### v3

@@ Abstract
The LogSumExp function, dual to the Kullback-Leibler (KL) divergence, plays a central role in many important optimization problems, including entropy-regularized optimal transport (OT) and distributionally robust optimization (DRO). In practice, when the number of exponential terms inside the logarithm is large or infinite, optimization becomes challenging since computing the gradient requires differentiating every term. We propose a novel convexity- and smoothness-preserving approximation to LogSumExp that can be efficiently optimized using stochastic gradient methods. This approximation is rooted in a sound modification of the KL divergence in the dual, resulting in a new $f$-divergence called the safe KL divergence. Our experiments and theoretical analysis of the LogSumExp-based stochastic optimization, arising in DRO and continuous OT, demonstrate the advantages of our approach over existing baselines.
log-sum-exp, stochastic optimization, distributionally robust optimization, optimal transport
@@ 1 Introduction
$F(\varphi;\mu)\coloneqq\log\int e^{\varphi(x)}\,\mathrm{d}\mu(x)\in(-\infty,\infty]$ (1)
@@
LogSumExp function appears commonly in optimization objectives, e.g., multiclass classification with softmax probabilities (Bishop and Nasrabadi, 2006), semi-dual formulation of entropy-regularized optimal transport (OT) (Peyré and Cuturi, 2019; Genevay et al., 2016), minimax problems (Pee and Royset, 2011), distributionally robust optimization (DRO) (Hu and Hong, 2012; Ben-Tal et al., 2013; Kuhn et al., 2025), maximum likelihood estimation (MLE) for exponential families and graphical models (Wainwright et al., 2008), variational Bayesian methods (Khan and Nielsen, 2018; Khan and Rue, 2023), information geometry (Amari and Nagaoka, 2000), KL-regularized Markov decision processes (Tiapkin et al., 2024). These problems involve minimizing $F(\varphi;\mu)$ w.r.t. a function $\varphi$, potentially parameterized by a vector $\theta$, e.g., a vector of neural network weights. Such optimization is characterized by two challenges. First, the decision variable $\varphi$ or $\theta$ often has large or infinite dimension. Second, the support of the measure $\mu$ can also be large or infinite. The first challenge is usually addressed by the use of first-order methods like Stochastic Gradient Descent (SGD), which are suitable for high-dimensional problems due to their cheap iterations. Unfortunately, there is no universal way to tackle the second challenge. If the number of exponential terms under the logarithm is large or infinite, i.e., the Monte Carlo estimate $\log\sum_{i=1}^{N}e^{\varphi(x_{i})}$ for a large $N$, the gradient computation requires differentiating each term, and, to the best of our knowledge, no cheap unbiased stochastic gradient have been proposed.
@@
In the current work, we propose a general-purpose approach to tackle both mentioned challenges. To that end, we use a SoftPlus approximation of $F(\varphi;\mu)$ that allows using stochastic gradient methods while remaining close to the original objective. We start with a variational formulation analogous to the one in the Gibbs principle, but with the KL-divergence replaced with another $f$-divergence – the safe KL divergence. The corresponding variational problem can be of interest itself, as it possesses some properties which can be beneficial compared to the KL penalty – e.g., uniform density bound. Moreover, it can be also viewed as an approximation of a conditional value at risk functional (CVaR). In fact, the same functional (with different parameters) appeared in (Soma and Yoshida, 2020) in the context of smooth CVaR approximation. Thus, we demonstrate that it generates a family of problems including CVaR and LogSumExp minimization as limit cases.
@@
Related works. We are not aware of any universal way to efficiently treat optimization objectives involving the LogSumExp functional (1), especially in infinite-dimensional settings. This functional appeared in different applications and was treated on a case-by-case basis. Bouchard (2007) studies three upper bounds on LogSumExp for approximate Bayesian inference. One of them is a particular case of the approximation proposed in the present work. (Titsias, 2016) constructs a bound on softmax probabilities and shows that it leads to a bound on LogSumExp in the context of multiclass classification. Nielsen and Sun (2016) approximate LogSumExp in the context of estimating divergences between mixture models. Their approximation combines LogSumExp bounds based on $\min$ and $\max$. Tucker et al. (2017); Luo et al. (2020) propose and study unbiased estimators for latent variable models based on Russian Roulette truncation. Lyne et al. (2015); Spring and Shrivastava (2017) focus on estimating the partition function itself, and do not consider questions of optimization involving the partition function. Hu and Hong (2012) and, subsequently, Levy et al. (2020) study DRO problems with $f$-divergences. They propose a batch-based approximation. When the ambiguity set is the unit simplex, and KL divergence penalty is used, the original objective is the LogSumExp of the losses over the entire dataset, while the approximation replaces it with the average of LogSumExp terms computed on individual batches. This approximation introduces bias, which can only be reduced by using large batch sizes. Deterministic LogSumExp maximization and minimization were considered in (Selvi et al., 2020) and (Kan et al., 2023), respectively. For stabilizing numerics related to evaluation of LogSumExp function, we refer to (Blanchard et al., 2021; Higham, 2021).
@@
    We empirically demonstrate the effectiveness of our approach on tasks, including computing continuous entropy-regularized OT and various DRO formulations. Our method outperforms existing state-of-the-art baselines in these applications and circumvents the overflow issue (Remark 3.1). It can also be combined with existing techniques. Therefore, it serves as a versatile tool for solving large-scale optimization problems.
@@
Notation. Given $a,a_{1},\dots,a_{n}\in\mathbb{R}$, we define
$\operatorname{LogSumExp}(a_{1},\dots,a_{n})\coloneqq\log\Bigl(\sum_{i=1}^{n}e^{a_{i}}\Bigr)$
and $\operatorname{SoftPlus}(a)\coloneqq\log(1+e^{a})$. Given a measurable space $\mathcal{X}$, by $\mathcal{P}(\mathcal{X})$ we denote the space of probability measures on $\mathcal{X}$, and by $\mathcal{C}(\mathcal{X})$ the space of continuous functions on $\mathcal{X}$. For $\mu,\nu\in\mathcal{P}(\mathcal{X})$ the Kullback–Leibler (KL) divergence is
@@ 2 SoftPlus approximation of log-partition function
$F(\varphi;\mu)=\sup_{\nu}\Big\{\int_{\mathcal{X}}\varphi(x)\,\mathrm{d}\nu(x)-D_{KL}(\nu,\mu):$
$\nu\in\mathcal{P}(\mathcal{X}),\,\int_{\mathcal{X}}|\varphi(x)|\,\mathrm{d}\nu(x)<\infty\Big\}$ (2)
with the maximum attained at the Gibbs measure $d\nu^{*}(x)=e^{\varphi(x)-F(\varphi;\mu)}d\mu(x),$ once $F(\varphi;\mu)<\infty$, see (Gibbs, 1902, Chapter XI, Theorem VI) or (Polyanskiy and Wu, 2025, Proposition 4.7) for the modern treatment.
@@
Clearly, $f_{\rho}(t)\to f_{0}(t)\coloneqq t\log t+1-t$ as $\rho\to 0$. Since $f_{0}$ induces the standard KL-divergence, $D_{\rho}$ is its approximation with accuracy regulated by the parameter $\rho$.
@@
$F_{\rho}(\varphi;\mu)\coloneqq\sup_{\nu}\left\{\int_{\mathcal{X}}\varphi(x)\,\mathrm{d}\nu(x)-D_{\rho}(\nu,\mu):\right.\\ \left.\nu\in\mathcal{P}(\mathcal{X}),\,\int_{\mathcal{X}}|\varphi(x)|\,\mathrm{d}\nu(x)<\infty\right\}.$ (5)
(i.e., $F_{\rho}(\cdot;\mu)$ is the convex conjugate of $D_{\rho}(\cdot,\mu)$). Note that the last term in $f_{\rho}$ prevents the density $\frac{d\nu}{d\mu}$ from being too large. In particular, it can not be greater than $\frac{1}{\rho}$. This can make the safe KL divergence a reasonable choice for unbalanced OT or DRO, as it imposes a hard constraint on the reweighting unlike the standard $D_{KL}$. Moreover, it can also be used instead of the entropy penalization in regularized OT (cf. capacity constrained transport in (Benamou et al., 2015, section 5.2)).
Again, by the convex duality and the variational principle (see Birrell et al., 2022, Theorem 6), we state the following properties.
@@
    If $\varphi(x)\leq M$ for all $x\in\mathcal{X}$, then
    $F_{\rho}(\varphi;\mu)\geq F(\varphi;\mu)-\rho e^{M-F(\varphi;\mu)}$
    for any $\rho\in\left(0,e^{F(\varphi;\mu)-M}\right)$.
@@
Let $a_{1},\dots,a_{n}\in\mathbb{R}$. Then for any $\rho\in(0,1)$
$\operatorname{LogSumExp}$ $(a_{1},\dots,a_{n})-\rho\leq$
$\leq\inf_{\alpha\in\mathbb{R}}\alpha-1+\frac{1}{\rho}\sum_{i=1}^{n}\log(1+\rho e^{a_{i}-\alpha})$
$\leq\operatorname{LogSumExp}(a_{1},\dots,a_{n}).$
For $\rho=1$ our approximation coincides with Bouchard’s bound for LogSumExp (Bouchard, 2007).
@@ 2.1 Links to CVaR
$\operatorname{CVaR}_{\rho}(\varphi;\mu)$ $\coloneqq\mathbb{E}_{X\sim\mu}\left[\varphi(X)\middle|\varphi(X)\geq Q_{1-\rho}\right]$
$=\frac{1}{\rho}\int_{\varphi(x)\geq Q_{1-\rho}}\varphi(x)\,\mathrm{d}\mu(x),$
where $Q_{1-\rho}$ is the $(1-\rho)$-quantile of $\varphi(X)$, $X\sim\mu$ (Rockafellar et al., 2000). Moreover, by Theorem 1 in (Rockafellar et al., 2000) CVaR also has the following formulation:
@@
Remarkably, in (Soma and Yoshida, 2020) the authors obtained a smooth approximation to CVaR which, up to an additive constant, has the same form as $F_{\rho}$. However, they considered the approximation w.r.t. a different parameter—a ”temperature” inside SoftPlus. Finally, (Levy et al., 2020) proposed another similar smoothed version of CVaR (KL-regularized CVaR) in the context of DRO. For our approximation, we obtain the following bounds.
@@
$\operatorname{CVaR}_{\rho}(\varphi;\mu)+\lambda(\log\rho-1)\leq\lambda F_{\rho}(\varphi/\lambda;\mu)\leq$
$\leq\operatorname{CVaR}_{\rho}(\varphi;\mu)+\lambda\left(\log\rho-1+\frac{1}{\rho}\right).$ (10)
@@ 2.2 The case of parametric models
In some applications, the function $\varphi$ is defined as the parametric loss function $L(x,\theta)$ and the goal is to minimize objective involving (1) w.r.t. $\theta$ to find the best model from the parametric family. In this section, we study our approximation to (1) in this parametric setting. Fix some closed parameter set $\Theta\subset\mathbb{R}^{d}$ and a loss function $L\colon\mathcal{X}\times\Theta\to\mathbb{R}$. Combining our approximation (6) and the minimization w.r.t. parameter $\theta$, we obtain the following minimization problem (note that we shifted $\alpha$ by $\log\rho$ compared to (6))
$\min_{\theta\in\Theta,\alpha\in\mathbb{R}}G_{\rho}(\theta,\alpha)\coloneqq\alpha+\log\rho-1+\\ +\frac{1}{\rho}\int_{\mathcal{X}}\log\left(1+e^{L(x,\theta)-\alpha}\right)\,\mathrm{d}\mu(x).$
@@
Note that
$f_{\rho}(t)=\frac{1}{\rho}\left((\rho t)\log(\rho t)+(1-\rho t)\log(1-\rho t)\right)+\\ +1-t\log\rho.$
Thus, unlike the KL entropy function $t\log t+1-t$, $f_{\rho}$ possesses the following favorable properties:
@@
The above properties are important from the computational optimization viewpoint. Recall that $\frac{d}{dt}\log(1+e^{t})=\frac{e^{t}}{1+e^{t}}\eqqcolon\sigma(t)$. Thus, we immediately obtain the following formulas for the gradient:
@@ 3 Applications
In this section we consider several particular applications involving the objective (1) and show numerically, that our general-purpose approach based on approximation (6) leads to better performance of SGD-type algorithms than the baseline algorithms designed specifically for these applications. The source code for all experiments is available at https://github.com/egorgladin/logsumexp-approx.
[…]